OpenStudy (anonymous):

Which statements are true? Choose all answers that are correct. A. 7 + (–10) < –13 + 17 B. 22 + (–13) > 18 + (–7) C. –18 + 24 = –23 + 17 D. –11 + (–8) > 4 + (–26)

OpenStudy (anonymous):

i need help with my studyguide

OpenStudy (anonymous):

fan and medal

OpenStudy (anonymous):

A because it is 4 and 4 is greater, D is correct because -18 is greater than -22

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

two more

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

What is the value of the variable that makes the statement true? a + (–7) + 3 = 0 A. –4 B. –3 C. 4 D. 7

OpenStudy (anonymous):

-7 + 3 = -4 If you take this step you will find your answer -4 + ? = 0

OpenStudy (anonymous):

What is the value of x + |y| when x = –11 and y = –4? A. –15 B. –7 C. 7 D. 15

OpenStudy (anonymous):

how do i give you a medal

OpenStudy (anonymous):

You already have.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

-11 + (-4 ) = -15

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Your welcome, if you want to give me another medal go to my profile page and click answered questions, and to answer your message sure I would love to be friends :)

OpenStudy (anonymous):

ok cool